Parth (parthkohli):

Why is the derivative of \(y^2\) with respect to \(x\) equal to \(2y \cdot y'\) or \(2y \cdot \large{dy \over dx}\)? I understand that the \(2y\) magic is power rule, but where does the \(y'\) come from? Just to add, I do know that \({dy \over dx} = y'\) or \(y'(x)\) :)

OpenStudy (anonymous):

this is the magic of chain rule !

Parth (parthkohli):

Right, can you remind me what it really said?

OpenStudy (anonymous):

derivative of y=x^2 is \[\frac{dy}{dx}=2x*\frac{d}{dx}(x)=2x*1=2x\] the same for y^2 produces \[\Large 2y*\frac{d}{dx}(y)=2y \frac{dy}{dx}\]

OpenStudy (anonymous):

implicit differentiation?

Parth (parthkohli):

Yes, Omni.

Parth (parthkohli):

I'm still trying to understand what @aliza_k said...

OpenStudy (anonymous):

ok what dont you understand

Parth (parthkohli):

I'm not really sure about the above.... wait a sec

OpenStudy (anonymous):

that is not the chain rule..

OpenStudy (anonymous):

power rule along with chain rule is \[y=x^n\] \[\frac{dy}{dx}=nx^{(n-1)}*\frac{d}{dx}(x)\]

OpenStudy (anonymous):

u need to differentiate the base as well.

Parth (parthkohli):

\[{d \over dx} f(g(x)) = f'(g(x))g'(x) \]

Parth (parthkohli):

That one then?

OpenStudy (anonymous):

Chain rule: \[\frac{dy}{dt} =\frac{dy}{dx} *\frac{dx}{dt} \] however, this is not required; only needed in related rates,

Parth (parthkohli):

How to apply this rule to\[{d \over dx} y^2\]

Parth (parthkohli):

Oh, got it... let me check.

OpenStudy (anonymous):

for the one you to wrote; chain rule is needed.

OpenStudy (anonymous):

\[\frac{d}{dx} y^2=\frac{d}{dy} \left(y^{2}\right) *\frac{dy}{dx} =2y*\frac{dy}{dx} \]

Parth (parthkohli):

Sorry... and yes! I understood the application :)

Parth (parthkohli):

\[{d \over dx}y^2 = 2y \times {dy \over dx}\]Yeah, so \(f\) is the square function and \(y(x)\) is the inner function :)

OpenStudy (anonymous):

yes

Parth (parthkohli):

\[y(x) = g\]

OpenStudy (anonymous):

i wouldnt use functions hard to understand :)

Parth (parthkohli):

... and then apply it to the rule I initially wrote

Parth (parthkohli):

Thanks people!

OpenStudy (anonymous):

thanks for your Question @ParthKohli Useful for me too ;)